write the equation of the line that is parallel to the line 4x+5y=7 and passes through the point (-1,8). thank you
Answer by lwsshak3(11628)

write the equation of the line that is parallel to the line 4x+5y=7 and passes through the point (-1,8). thank you
**
Standard form of equation for a straight line: y=mx+b, m= slope, b=y-intercept.
..
given equation: 4x+5y=7
rewrite in standard form
5y=-4x+7
y=-4x/5+7/5
m=-4/5
..
Parallel equation has same slope=-4/5
y=-4x/5+b
solving for b using coordinates of given point thru which the line passes.
8=-4(-1)/5+b
8=4/5+b
b=8-4/5=40/5-4/5=36/5
Equation: y=-4x/5+36/5